OpenStudy (anonymous):

How do I identify the intercepts, removable continuities, and asymptotes? (2x^3-3x+1)/x^3-5x+7

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

are you allowed to graph the function ?

OpenStudy (anonymous):

nope :/ i think i'm supposed to solve it based on just the equation

ganeshie8 (ganeshie8):

Finding horizontal asymptotes is easy for a rational function

ganeshie8 (ganeshie8):

\[y = \dfrac{\color{Red}{2}x^3-3x+1}{\color{red}{1}x^3-5x+7}\] As \(x\) gets large, the lower degree terms don't matter much, so the value approaches \[y \approx \dfrac{\color{Red}{2}x^3}{\color{red}{1}x^3} = \dfrac{\color{Red}{2}}{\color{red}{1}}=\color{red}{2}\] so \(\color{red}{y=2}\) is an horizontal asymptote

OpenStudy (anonymous):

thank you!! is the vertical asymptote about 2.75?

OpenStudy (anonymous):

x^3-5x=-7 right? and solve for x

ganeshie8 (ganeshie8):

it should be around x= -2.75 but how did u get that ?

ganeshie8 (ganeshie8):

Yes

OpenStudy (anonymous):

okay thanks!! i'm super confused on finding the x intercepts though

ganeshie8 (ganeshie8):

did you find y intercepts ?

OpenStudy (anonymous):

yeah i got 1/7 by plugging in 0 in place of x

ganeshie8 (ganeshie8):

Good. use the same trick simply plugin y=0 to get the x intercepts

ganeshie8 (ganeshie8):

\[y = \dfrac{\color{black}{2}x^3-3x+1}{\color{black}{1}x^3-5x+7}\] plugin y=0 for x intercepts \[0 = \dfrac{\color{black}{2}x^3-3x+1}{\color{black}{1}x^3-5x+7}\]

ganeshie8 (ganeshie8):

which is same as \[0 =\color{black}{2}x^3-3x+1\] solve

ganeshie8 (ganeshie8):

wolfram gives 3 x intercepts http://www.wolframalpha.com/input/?i=solve+2x%5E3-3x%2B1%3D0

OpenStudy (anonymous):

could i use the formula bc of the plus and minus?

OpenStudy (anonymous):

quadratic formula*

ganeshie8 (ganeshie8):

nope, the equation is cubic, not quadratic. i would simply use calculator/wolfram

OpenStudy (anonymous):

i'm confused on how i would get 3 answers :/

ganeshie8 (ganeshie8):

why ? it is fine to have multiple x intercepts

ganeshie8 (ganeshie8):

that just means, the graph cuts the x axis at 3 different places

OpenStudy (anonymous):

i used mathway and it gave me (1,0) and (-1+-sqrt3)/2 but i'm not sure how it got that

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

oh i see it haha thank you so much :)